# CBAM-EfficientNetV2 for Histopathology Image Classification using Transfer Learning and Dual Attention Mechanisms

### v5

@@ Abstract
Breast cancer histopathology image classification is critical for early detection and improved patient outcomes. This study introduces a novel approach leveraging EfficientNetV2 models, to improve feature extraction and focus on relevant tissue regions. The proposed models were evaluated on the BreakHis dataset across multiple magnification scales (40X, 100X, 200X, and 400X). Among them, the EfficientNetV2-XL with CBAM achieved outstanding performance, reaching a peak accuracy of 98.96% and an F1-score of 98.31% at 400X magnification, outperforming state-of-the-art methods. By integrating Contrast Limited Adaptive Histogram Equalization (CLAHE) for preprocessing and optimizing computational efficiency, this method demonstrates its suitability for real-time clinical deployment. The results underscore the potential of attention-enhanced scalable architectures in advancing diagnostic precision for breast cancer detection.
@@
Transfer Learning, Modality Specific, EfficientNet, CBAM, Histopathology, Image Classification, Deep Learning, Breast Cancer.
@@ I Introduction
Automating histopathological analysis has become a vital research area in medical imaging due to its potential to significantly reduce the diagnostic workload of pathologists and enhance the consistency and reproducibility of results. Histopathology—the microscopic study of tissue abnormalities—is fundamental in diagnosing diseases such as cancer, but manual analysis is labor-intensive, subjective, and prone to variability between observers. As a result, computer-aided diagnosis (CAD) systems have gained traction, with the aim of augmenting pathologists’ abilities to efficiently and accurately classify tissue samples [1]. Despite substantial progress, CAD systems continue to face significant challenges, particularly in handling the variability inherent in histopathological images. Differences in magnification levels, staining protocols (e.g., Hematoxylin and Eosin, or H&E staining), and tissue morphology introduce variability that complicates the generalization of standard deep learning models [2].
Traditional convolutional neural networks (CNNs), while successful in many computer vision tasks, often struggle to generalize across these variations in histopathological data. Specifically, CNNs are designed with fixed receptive fields and are limited in their ability to capture fine-grained details that vary across multiple scales [3]. The heterogeneity of breast cancer histopathology images—marked by different magnification levels and complex cellular structures—poses unique challenges that demand more sophisticated architectures. To address this, researchers have turned to hybrid models that integrate attention mechanisms with CNNs, which have demonstrated superior performance in medical imaging tasks by enhancing feature extraction.
Attention mechanisms such as the Convolutional Block Attention Module (CBAM) have been adapted to improve image classification by enabling the network to focus dynamically on the most informative regions of an image. In histopathology, these mechanisms are particularly effective in highlighting regions critical to cancer diagnosis, such as irregular nuclei shapes, abnormal cellular arrangements, and other key tissue structures that may indicate malignancy. By applying attention, the network can dynamically recalibrate feature maps along spatial and channel dimensions, improving its ability to capture both local and global image features [4]. This is especially valuable in distinguishing between benign and malignant tissue, where small but significant differences in morphology must be detected.
EfficientNet [8], a state-of-the-art CNN architecture known for its compound scaling strategy, is particularly well-suited for medical image classification tasks due to its ability to balance model depth, width, and resolution. EfficientNet achieves superior performance with fewer parameters, making it computationally efficient without sacrificing accuracy. By integrating attention mechanisms with EfficientNet, we aim to build models that can more effectively classify breast cancer histopathology images while remaining scalable to real-world clinical applications. This hybrid approach addresses the need for models that can operate across varying image resolutions and magnification levels, ensuring robustness and high accuracy in diverse settings.
@@ III-E Training and Hyperparameter Tuning
TABLE I: Performance of Base Models Across Magnification Scales
Model Scale Accuracy (%) F1 Score (%)
EfficientNetV2-XL (Base) 40X 86.78 84.23
100X 87.89 85.34
200X 88.23 86.67
400X 89.12 87.56
EfficientNetV2-L (Base) 40X 85.67 83.13
100X 86.89 84.23
200X 87.13 85.56
400X 87.89 86.23
@@ IV-A Results
We employed a modality-specific transfer learning strategy for the classification task and compared it with various state-of-the-art CNN architectures, as shown in Table II.
@@
TABLE II: Comparison of Accuracy Across Models and Magnification Factors
@@
Proposed Hybrid 2024 40X 96.11
(EfficientNetV2-XL + CBAM) 100X 98.04
@@ V Conclusion
CBAM-EfficientNetV2-XL achieved superior accuracy of 99.01%, however EfficientNetv2-XL has F1-Score of 98.31 ,therefore EfficientNetV2-XL + CBAM outperforms other models in accuracy. By leveraging attention mechanisms, the model effectively focuses on critical regions of histopathological images, making it well-suited for real-time clinical diagnostics. These results demonstrate that the proposed model surpasses the performance of previous state-of-the-art models, highlighting its potential for real-time clinical deployment. Future work may adapt the model to other datasets and further expand its clinical applicability.

### v6

# Breast Cancer Histopathology Classification using CBAM-EfficientNetV2 with Transfer Learning

@@ Abstract
Breast cancer histopathology image classification is crucial for early diagnosis, enabling timely interventions and improved patient outcomes. This study proposes a novel CBAM-EfficientNetV2 model that integrates Convolutional Block Attention Modules (CBAM) with EfficientNetV2 to enhance feature extraction and focus on clinically relevant tissue regions. Leveraging transfer learning and attention mechanisms, the model achieves superior performance on the BreakHis dataset across multiple magnifications (40X, 100X, 200X, 400X), with an accuracy of 99.01% and an F1-score of 98.31% at 400X, outperforming state-of-the-art methods. Robust preprocessing with Contrast Limited Adaptive Histogram Equalization (CLAHE) and optimized computational efficiency ensure scalability for real-time clinical diagnostics. This approach advances responsible healthcare innovation, promoting accessible, high-precision breast cancer detection for global health equity.
@@
Transfer Learning, EfficientNetV2, CBAM, Histopathology, Breast Cancer Detection, Deep Learning, Medical Image Classification, Responsible Healthcare Technology
@@ I Introduction
Automating histopathological analysis is transforming medical imaging by reducing diagnostic workloads and enhancing the accuracy and reproducibility of disease detection. Histopathology, the microscopic examination of tissue abnormalities, is critical for diagnosing conditions like breast cancer, yet manual analysis remains labor-intensive, subjective, and prone to inter-observer variability [1]. Computer-aided diagnosis (CAD) systems address these challenges by augmenting pathologists’ capabilities, offering scalable solutions for consistent and efficient tissue classification [2]. However, the variability in histopathological images—due to differences in magnification levels, staining protocols (e.g., Hematoxylin and Eosin, H&E), and tissue morphology—poses significant hurdles for standard deep learning models, limiting their generalization across diverse clinical scenarios [3].
Conventional convolutional neural networks (CNNs), while effective in general computer vision, often struggle to capture the multi-scale, fine-grained details inherent in breast cancer histopathology images [3]. Fixed receptive fields in CNNs restrict their ability to adapt to the heterogeneous nature of these images, which vary across magnifications (40X to 400X) and exhibit complex cellular structures. To overcome these limitations, hybrid architectures combining CNNs with attention mechanisms have emerged, enabling models to focus on clinically relevant features and improve diagnostic precision.
This study leverages EfficientNetV2 [8], a state-of-the-art CNN known for its compound scaling strategy, which optimizes depth, width, and resolution for superior performance with fewer parameters. By integrating Convolutional Block Attention Modules (CBAM), we enhance EfficientNetV2’s ability to prioritize critical tissue regions, addressing the variability in histopathological data. This hybrid approach delivers high accuracy and computational efficiency, making it ideal for real-time clinical deployment. Aligned with the vision of responsible innovation, our work promotes sustainable healthcare solutions by enabling precise, accessible breast cancer diagnostics across diverse clinical settings.
@@ IV-A Results
We employed a modality-specific transfer learning strategy for the classification task and compared it with various state-of-the-art CNN architectures, as shown in Table I.
@@
TABLE I: Comparison of Accuracy Across Models and Magnification Factors
@@
EfficientNetV2-XL + CBAM CBAM-EfficientNetV2 40X 96.11
100X 98.04
@@ V Conclusion
CBAM-EfficientNetV2-XL achieved superior accuracy of 99.01%, therefore EfficientNetV2-XL + CBAM outperforms other models in accuracy. By leveraging attention mechanisms, the model effectively focuses on critical regions of histopathological images, making it well-suited for real-time clinical diagnostics. These results demonstrate that the proposed model surpasses the performance of previous state-of-the-art models. Future work may adapt the model to other datasets and further expand its clinical applicability.